# Theory of local $\mathbb{Z}_{2}$ topological markers for finite and periodic two-dimensional systems

### v1

@@ III.1 Local OBC spin-Chern marker
Consider a Hamiltonian $\mathcal{H}$ for which $[\mathcal{H},S_{z}]=0$, where $S_{z}$ is the spin operator along the $z$-direction. Since $\mathcal{H}$ commutes with $S_{z}$, it can be decoupled into spin sectors that are mapped onto each other by TR symmetry. Such splitting allows the definition of two individual Chern numbers $C_{\uparrow,\downarrow}$, physically related to the quantized spin Hall conductivity [47], from which the $\mathbb{Z}_{2}$ invariant of the system can be computed as:
@@
where the two individual Chern numbers can be obtained by integrating the Berry curvature of the relative spin-up or spin-down subspace. The states composing the two spin sectors can be selected by computing the ground-state projector $\mathcal{P}(\mathbf{k})$ and diagonalizing $\mathcal{P}(\mathbf{k})S_{z}\mathcal{P}(\mathbf{k})$, whose spectrum has eigenvalues ±$\pm\frac{1}{2}$. Remarkably, Prodan [37] proved that this strategy holds even if $[\mathcal{H},S_{z}]\neq 0$ as long as the operator $\mathcal{P}(\mathbf{k})S_{z}\mathcal{P}(\mathbf{k})$ displays a gap in its spectrum. In this case, if one diagonalize $\mathcal{P}(\mathbf{k})S_{z}\mathcal{P}(\mathbf{k})$, its eigenvalues $s_{\lambda}$ will spread symmetrically about zero in the interval $[-\frac{1}{2},\frac{1}{2}]$, and the projectors onto the positive ($\sigma=+$) and negative ($\sigma=-$) eigenvalues can be defined:
$\mathcal{P}_{\sigma}(\mathbf{k})=\sum_{\lambda:\mathrm{sgn}(s_{\lambda})=\sigma}\ket{\phi_{\lambda\mathbf{k}}}\bra{\phi_{\lambda\mathbf{k}}}$ | (10)
@@
so, due to TR symmetry, it must hold that $C_{+}=-C_{-}$, and the $\mathbb{Z}_{2}$ invariant can be computed as a spin-Chern number, as in Eq. 9, by substituting $C_{\uparrow}\rightarrow C_{+}$ and $C_{\downarrow}\rightarrow C_{-}$.
Within OBCs the BZ does not exist, and to obtain a local spin-Chern marker (LSCM) it is possible to apply the LCM formula [Eq. 3] to each of the individual Chern numbers appearing in Eq. 9. By doing so, two individual local Chern markers can be computed:
@@
whose macroscopic average is a local $\mathbb{Z}_{2}$ invariant. Being expressed as a trace, the LSCM can be evaluated in real space as for the OBC LCM. We stress again that such formulation is useful only within OBCs, as the general way to calculate the operators $[r_{\alpha},\mathcal{P}_{\sigma}]$ requires the explicit evaluation of the terms $r_{\alpha}\mathcal{P}_{\sigma}$ and $\mathcal{P}_{\sigma}r_{\alpha}$, which are ill-defined within PBCs. We will discuss how to generalize these local markers to the PBC framework in Sec. IV. As stated before, the LSCM remains well-defined whenever the projected spin operator displays a spectral gap, meaning that it can also be computed for systems breaking TR symmetry. However, in general, the existence of a gap in $\mathcal{P}S_{z}\mathcal{P}$ is not guaranteed, as it is not a fundamental property of QSHIs. Hence, the LSCM is a valid approach but not, in principle, a truly general formulation of the $\mathbb{Z}_{2}$ invariant.
@@ III.2 Local OBC $\mathbb{Z}_{2}$ marker based on time-reversal symmetry
QSHIs are characterized by the TR symmetry $\Theta$, which forces the Chern number of the system to vanish, and the classification of topological phases to be $\mathbb{Z}_{2}$, resulting in a symmetry-protected topological phase. Hence, it seems reasonable that a general definition of the $\mathbb{Z}_{2}$ topological invariant should exploit TR symmetry explicitly, similarly to what has been done, for instance, for the mirror-Chern number [46, 48]. However, the TR operator $\Theta$ is antiunitary and it cannot be diagonalized to define a topological invariant. Nonetheless, we can still split the occupied states into two Chern subspaces using $\Theta$ thanks to Kramers’ theorem. A similar strategy was proven to be successful in computing the global invariant of a homogeneous and periodic QSHI with only two occupied bands [39]. The calculation of the topological invariant, in this case, is achieved by performing a careful parallel transport procedure in almost all the BZ: If one insists in having a periodic and TR-symmetric gauge, this cannot be smooth due to the topological obstruction arising in the topological phase. To evaluate the $\mathbb{Z}_{2}$ invariant one can impose a TR symmetric gauge that is smooth on the BZ cylinder (but not on the BZ torus, i.e., a non-periodic gauge), disentangling the occupied bands into two Chern manifolds, each characterized by its own integer individual Chern number [39].
@@
Within OBCs, the parallel transport procedure becomes irrelevant, and the only meaningful step is the TR constraint. Numerical diagonalizations carry random global phases for each eigenvector, so to impose a TR symmetric gauge we need to select an initial state in the degenerate subspace and explicitly evaluate its TR partner. If the dimension of the degenerate subspace is greater than two, then an orthonormalization procedure of the remaining eigenvectors with respect to the ones considered before is needed in order to get a set of orthonormal basis vectors related by TR symmetry. This procedure can be repeated until the number of vectors is equal to the dimension of the degenerate subspace. By doing so, the eigenstates of the Hamiltonian are split in two orthogonal subsets (identified by the projectors $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$) that individually break TR, and that can be mapped onto each other by symmetry. That allows us to evaluate the individual local Chern numbers of the two subsets, and obtain a local $\mathbb{Z}_{2}$ marker. However, this construction in general does not result in exponentially localized projectors $\mathcal{P}_{1,2}$, and their decay properties in the bulk depend on the choice of the eigenstates assigned to each subspace.
@@
One way to obtain exponentially localized projectors is to build the maximally localized Wannier functions (MLWF) of the system [49, 50] (also known as Boys orbitals in OBCs [51]). TR symmetric systems are characterized by a vanishing Chern number, which in principle should allow computing exponentially localized Wannier functions (WF). In practice, however, the usual method to compute them is via a projection procedure [50], which is ensured to fail in the topological phase if the trial states are TR symmetric [52]. This is a consequence of the topological obstruction, that forbids the existence of a smooth and TR-symmetric gauge in the topological phase [39]. Within OBCs then, once a set of trial functions $g_{n}(\mathbf{r})$ is chosen, these are projected onto the occupied orbitals:
|$\ket{\phi_{n}}=\sum_{m=1}^{J}\ket{\psi_{n}}\braket{\psi_{n}}{g_{n}}.$ (14)
@@
MLWFs can then be split by using the TR symmetry as explained above. This procedure allows us to define a local $\mathbb{Z}_{2}$ marker (LZ2M) that is uniquely based on TR symmetry as
@@
where $\mathfrak{C}_{1}(\mathbf{r})$ and $\mathfrak{C}_{2}(\mathbf{r})$ are computed via Eq. 3 using the projectors $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ respectively. It is worth noting that, being the LZ2M based entirely on TR symmetry, its formulation does not rely on the existence of a spectral gap in any operator (but the Hamiltonian), thus making it a more general formulation of the local $\mathbb{Z}_{2}$ marker with respect to the LSCM. However, the fact that the LZ2M returns the correct value of the marker strongly depends on the choice of the trial functions |$\ket{g_{n}}$ introduced in Eq. 14. Specifically, as explained in Sec. VI.1, in order to avoid the topological obstruction, the trial functions |$\ket{g_{n}}$ should not be related by TR symmetry. This condition, in turn, implies that $\mathcal{P}\neq\mathcal{P}_{1}+\Theta\mathcal{P}_{1}\Theta^{-1}$. However, since we want to split the occupied Hilbert space into two subsets each representative for the whole space and related by TR symmetry (to compute the $\mathbb{Z}_{2}$ invariant from Chern numbers), we need the two projectors $\mathcal{P}$ and $\mathcal{P}_{\Theta}=\mathcal{P}_{1}+\Theta\mathcal{P}_{1}\Theta^{-1}$ to be as close as possible [10]. For this reason, we require that the choice of the trial projection, discussed in Sec. VI.1, should also minimize the spillage between the projectors $\mathcal{P}$ and $\mathcal{P}_{\Theta}$, defined as 111 Since the trace of $\mathcal{P}$ is equal to the number of occupied states $N_{occ}$, this formulation of the spillage is equivalent to the more common expression in terms of the projector onto empty states: $\mathrm{Tr}[(\mathcal{P}-\mathcal{P}_{\Theta})^{2}]/(2N_{occ})=\mathrm{Tr}[\mathcal{P}\mathcal{Q}_{\Theta}]/N_{occ}=\mathrm{Tr}[\mathcal{Q}\mathcal{P}_{\Theta}]/N_{occ}$.:
@@ IV Local $\mathbb{Z}_{2}$ markers for periodic systems
IV.1 Local PBC spin-Chern number
Within PBCs, following the same strategy developed in Sec. III.1, the individual Chern numbers $C_{\pm}$ introduced in Eq. 9 can be promoted to local individual Chern markers using the PBC LCM [Eq. 7]. A complementary, and equivalent, way to look at this strategy is that we derive from single-point spin-Chern numbers introduced in Ref. [12] the corresponding local markers. In the limit of a very large supercell, the BZ shrinks to the $\Gamma$-point only, so just one diagonalization of the Hamiltonian is required. Defining $\mathcal{P}_{\Gamma}=\sum_{n=1}^{N_{occ}}\ket{u_{n\Gamma}}\bra{u_{n\Gamma}}$ | the ground-state projector, we can diagonalize the projected spin operator $\left(\mathcal{P}_{\Gamma}S_{z}\mathcal{P}_{\Gamma}\right)\ket{\phi_{\lambda\Gamma}}=s_{\lambda}\ket{\phi_{\lambda\Gamma}}$, and construct the dual states for each the spin sector:
|$\ket{\tilde{\phi}_{\lambda\mathbf{b}_{j}}}=\sum_{\mu:\mathrm{sgn}(s_{\mu})=\sigma}S_{\mu\lambda}^{-1}(\mathbf{b}_{j})e^{-i\mathbf{b}_{j}\cdot\mathbf{r}}\ket{\phi_{\mu\Gamma}}$ (19)
@@ IV.1 Local PBC spin-Chern number
$\mathcal{P}_{\mathbf{b}_{j}}^{\pm}=\sum_{\lambda:\mathrm{sgn}(s_{\lambda})=\pm}\ket{\tilde{\phi}_{\lambda\mathbf{b}_{j}}}\bra{\tilde{\phi}_{\lambda\mathbf{b}_{j}}},$ (20)
@@
where $\mathcal{P}_{\Gamma}^{\pm}=\sum_{\lambda:\mathrm{sgn}(s_{\lambda})=\pm}\ket{\phi_{\lambda\Gamma}}\bra{\phi_{\lambda\Gamma}}$ |, so that it holds that $\mathcal{P}_{\Gamma}=\mathcal{P}_{\Gamma}^{+}+\mathcal{P}_{\Gamma}^{-}$. In particular, the covariant derivative is approximated by the forward finite difference formula in Eq. 21 and by the symmetric finite difference in Eq. 22. Finally, the PBC LSCM can be defined as:
@@
IV.2 Local PBC $\mathbb{Z}_{2}$ markers based on time-reversal symmetry
@@ V.1 Kane-Mele model
where $c_{i}^{\dagger}=(c_{i\uparrow}^{\dagger},c_{i\downarrow}^{\dagger})$ and sums on spin indices are implied, with the convention that if no spin matrices appear, they are contracted over the identity. In Equation V.1, $t$ is the nearest-neighbor hopping amplitude, and $\Delta$ is a staggered on-site potential, depending on the sublattice identified by $\tau_{i}\in\{0,1\}$. In the following, we will set $t=1$. The parameter $\lambda_{SO}$ is the intensity of the diagonal spin-orbit coupling, introduced as a complex hopping between second nearest-neighbors, where $\nu_{ij}=\mathrm{sign}(\mathbf{d}_{1}\times\mathbf{d}_{2})_{z}$ accounts for the direction of the hopping and $\mathbf{d}_{1,2}$ are unit vectors connecting second nearest-neighbors. Finally, $\lambda_{R}$ is the amplitude of the Rashba term, which couples the two spin sectors and breaks $S_{z}$ symmetry. Here $\boldsymbol{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ is the vector of Pauli matrices, and ${\mathbf{e}}_{\langle ij\rangle}=\mathbf{d}_{\langle ij\rangle}\times\hat{\mathbf{z}}$ where $\mathbf{d}_{\langle ij\rangle}$ is the unit vector along the direction connecting site $i$ to site $j$.

### v2

@@ III.1 Local OBC spin-Chern marker
We start by considering a Hamiltonian $\mathcal{H}$ for which $[\mathcal{H},S_{z}]=0$, where $S_{z}$ is the spin operator along the $z$-direction. Since $\mathcal{H}$ commutes with $S_{z}$, it can be decoupled into spin sectors that are mapped onto each other by TR symmetry. Such splitting allows the definition of two individual Chern numbers $C_{\uparrow,\downarrow}$, physically related to the quantized spin Hall conductivity [53], from which the $\mathbb{Z}_{2}$ invariant of the system can be computed as:
@@
where the two individual Chern numbers can be obtained by integrating the Berry curvature of the relative spin-up or spin-down subspace. The states composing the two spin sectors can be selected by computing the ground-state projector $\mathcal{P}(\mathbf{k})$ and diagonalizing $\mathcal{P}(\mathbf{k})S_{z}\mathcal{P}(\mathbf{k})$, whose spectrum has eigenvalues ±$\pm\frac{1}{2}$. Remarkably, Prodan [42] proved that this strategy holds even if $[\mathcal{H},S_{z}]\neq 0$ as long as the operator $\mathcal{P}(\mathbf{k})S_{z}\mathcal{P}(\mathbf{k})$ displays a gap in its spectrum. In this case, if one diagonalize $\mathcal{P}(\mathbf{k})S_{z}\mathcal{P}(\mathbf{k})$, its eigenvalues $s_{\lambda}$ will spread symmetrically around zero in the interval $[-\frac{1}{2},\frac{1}{2}]$, and the projectors onto the positive ($\sigma=+$) and negative ($\sigma=-$) eigenvalues can be defined:
$\mathcal{P}_{\sigma}(\mathbf{k})=\sum_{\lambda:\mathrm{sign}(s_{\lambda})=\sigma}\ket{\phi_{\lambda\mathbf{k}}}\bra{\phi_{\lambda\mathbf{k}}},$ (10)
@@
so, due to TR symmetry, it must hold that $C_{+}=-C_{-}$, and the $\mathbb{Z}_{2}$ invariant can be computed as a spin-Chern number, as in Eq. 9, by substituting $C_{\uparrow}\rightarrow C_{+}$ and $C_{\downarrow}\rightarrow C_{-}$. It should be noted that other definitions of the spin-Chern number exists, for instance as the integral of the spin-Berry curvature [30]. However, while Prodan’s spin-Chern number [42] is quantized, the integral of the spin-Berry curvature is only an approximate indicator of the topology of the system, although experimentally relevant since related to the spin-resolved local circular dichroism [30].
Within OBCs the BZ does not exist: A local spin-Chern marker (LSCM) can then be obtained by applying the LCM formula [Eq. 3] to each of the individual Chern numbers appearing in Eq. 9. By doing so, two individual local Chern markers can be computed:
@@
whose macroscopic average is a local $\mathbb{Z}_{2}$ invariant. Being expressed as a trace, the LSCM can be evaluated in real space as for the OBC LCM. A similar strategy can also be used to define a layer-resolved spin-Chern marker in 3D systems [33]. We stress again that such formulation is useful mostly within OBCs, as the general way to calculate the operators $[r_{\alpha},\mathcal{P}_{\sigma}]$ requires the explicit evaluation of the terms $r_{\alpha}\mathcal{P}_{\sigma}$ and $\mathcal{P}_{\sigma}r_{\alpha}$, which are ill-defined within PBCs. We will discuss how to generalize these local markers to the PBC framework in Sec. IV. As stated before, the LSCM remains well-defined whenever the projected spin operator displays a spectral gap, possibly also for systems breaking TR symmetry. However, in general, the existence of a gap in $\mathcal{P}S_{z}\mathcal{P}$ is not guaranteed, as it is not a fundamental property of QSHIs. Hence, the LSCM is a valid approach but not, in principle, a truly general formulation of the $\mathbb{Z}_{2}$ invariant.
@@ III.2 Local OBC $\mathbb{Z}_{2}$ marker based on time-reversal symmetry
QSHIs are characterized by the TR symmetry $\Theta$, which forces the Chern number of the system to vanish, and the classification of topological phases to be $\mathbb{Z}_{2}$, resulting in a symmetry-protected topological phase. Hence, it seems reasonable that a general definition of the $\mathbb{Z}_{2}$ topological invariant should exploit TR symmetry explicitly, similarly to what has been done, for instance, for the mirror-Chern number [52, 54]. However, the TR operator $\Theta$ is antiunitary and it cannot be diagonalized to define a topological invariant. Nonetheless, we can still split the occupied states into two Chern subspaces using $\Theta$ thanks to Kramers’ theorem. A similar strategy was proven to be successful in computing the global invariant of a homogeneous and periodic QSHI with only two occupied bands [44]. The calculation of the topological invariant, in this case, is achieved by enforcing the parallel-transport gauge in almost all the BZ: If one insists in having a periodic and TR-symmetric gauge, this cannot be smooth due to the topological obstruction arising in the topological phase. To evaluate the $\mathbb{Z}_{2}$ invariant one can impose a TR symmetric gauge that is smooth on the BZ cylinder (but not on the BZ torus, i.e., a non-periodic gauge), disentangling the occupied bands into two Chern manifolds, each characterized by its own integer individual Chern number [44].
@@
Within OBCs, the parallel transport procedure becomes irrelevant, and the only meaningful step is the TR constraint. Numerical diagonalizations carry random global phases for each eigenvector, so to impose a TR symmetric gauge we need to select an initial state in the degenerate subspace and explicitly evaluate its TR partner. If the dimension of the degenerate subspace is greater than two, then an orthonormalization procedure is needed. This procedure can be repeated until the number of vectors is equal to the dimension of the degenerate subspace. By doing so, the eigenstates of the Hamiltonian are split in two orthogonal subsets (identified by the projectors $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$) that individually break TR, and that can be mapped onto each other by symmetry. That allows us to evaluate the individual local Chern numbers of the two subsets, and obtain a local $\mathbb{Z}_{2}$ marker. However, this construction in general does not result in exponentially localized projectors $\mathcal{P}_{1,2}$, and their decay properties in the bulk depend on the choice of the eigenstates assigned to each subspace.
@@
One way to obtain exponentially localized projectors is to build the maximally localized Wannier functions (MLWF) of the system [55, 56] (also known as Boys orbitals in OBCs [57]). TR symmetric systems are characterized by a vanishing Chern number, which in principle should allow computing exponentially localized Wannier functions (WF). In practice, however, the simplest method to compute them is via a projection procedure [56], which is ensured to fail in the topological phase if the trial states are TR symmetric [58]. This is a consequence of the topological obstruction, that forbids the existence of a smooth and TR-symmetric gauge in the topological phase [44]. Within OBCs, a set of trial functions $g_{n}(\mathbf{r})$ can be chosen based on chemical intuition. Then, these are used to project the $J$ occupied orbitals:
|$\ket{\phi_{n}}=\sum_{m=1}^{J}\ket{\psi_{m}}\braket{\psi_{m}}{g_{n}}.$ (14)
@@
MLWFs can then be split in two sets by using TR symmetry as explained above. This procedure allows us to define a local $\mathbb{Z}_{2}$ marker (LZ2M) that is uniquely based on TR symmetry as
@@
where $\mathfrak{C}_{1}(\mathbf{r})$ and $\mathfrak{C}_{2}(\mathbf{r})$ are computed via Eq. 3 with the projectors $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ respectively. It is worth noting that, as the LZ2M is based entirely on TR symmetry, its formulation does not rely on the existence of a spectral gap in any operator (but the Hamiltonian), thus being more general with respect to the LSCM. However, the validity of the LZ2M strongly depends on the choice of the trial functions |$\ket{g_{n}}$ introduced in Eq. 14. Specifically, as explained in Sec. VI.1, in order to avoid the topological obstruction, the trial functions |$\ket{g_{n}}$ should not be related by TR symmetry. This condition, in turn, implies that $\mathcal{P}\neq\mathcal{P}_{1}+\Theta\mathcal{P}_{1}\Theta^{-1}$. However, since we want to split the occupied Hilbert space into two subsets that represents the whole space and are related by TR symmetry (to compute the $\mathbb{Z}_{2}$ invariant from Chern numbers), we need the two projectors $\mathcal{P}$ and $\mathcal{P}_{\Theta}=\mathcal{P}_{1}+\Theta\mathcal{P}_{1}\Theta^{-1}$ to be as close as possible [10]. For this reason, we require that the choice of the trial projection, discussed in Sec. VI.1, should also minimize the spillage between the projectors $\mathcal{P}$ and $\mathcal{P}_{\Theta}$, defined as 111 The spillage can be interpreted as a measure of how much two projectors describe the same subspace, or equivalently, a distance between manifolds. Since the trace of $\mathcal{P}$ is equal to the number of occupied states $N_{occ}$, this formulation of the spillage is equivalent to the more common expression in terms of the projector onto empty states: $\mathrm{Tr}[(\mathcal{P}-\mathcal{P}_{\Theta})^{2}]/(2N_{occ})=\mathrm{Tr}[\mathcal{P}\mathcal{Q}_{\Theta}]/N_{occ}=\mathrm{Tr}[\mathcal{Q}\mathcal{P}_{\Theta}]/N_{occ}$. From these forms it is easier to see that, for $\gamma=0$, the projectors describe the same subspace, while if $\gamma=1$, the two projectors are orthogonal and thus describe orthogonal subspaces.:
@@ IV Local $\mathbb{Z}_{2}$ markers for periodic systems
IV.1 Local PBC spin-Chern marker
Within PBCs, following the same strategy developed in Sec. III.1, the individual Chern numbers $C_{\pm}$ introduced in Eq. 9 can be promoted to local individual Chern markers using the PBC LCM [Eq. 7]. A complementary, and equivalent, way to look at this strategy is that we derive from single-point spin-Chern numbers introduced in Ref. [12] the corresponding local markers. In the limit of a very large supercell, the BZ shrinks to the $\Gamma$-point only, so just one diagonalization of the Hamiltonian is required. Defining $\mathcal{P}_{\Gamma}=\sum_{n=1}^{N_{occ}}\ket{u_{n\Gamma}}\bra{u_{n\Gamma}}$ | the ground-state projector, we can diagonalize the projected spin operator $\left(\mathcal{P}_{\Gamma}S_{z}\mathcal{P}_{\Gamma}\right)\ket{\phi_{\lambda\Gamma}}=s_{\lambda}\ket{\phi_{\lambda\Gamma}}$, and construct the dual states for each spin sector:
|$\ket{\tilde{\phi}_{\lambda\mathbf{b}_{j}}}=\sum_{\mu:\mathrm{sign}(s_{\mu})=\sigma}S_{\mu\lambda}^{-1}(\mathbf{b}_{j})e^{-i\mathbf{b}_{j}\cdot\mathbf{r}}\ket{\phi_{\mu\Gamma}}$ (19)
@@ IV.1 Local PBC spin-Chern marker
$\mathcal{P}_{\mathbf{b}_{j}}^{\pm}=\sum_{\lambda:\mathrm{sign}(s_{\lambda})=\pm}\ket{\tilde{\phi}_{\lambda\mathbf{b}_{j}}}\bra{\tilde{\phi}_{\lambda\mathbf{b}_{j}}},$ (20)
@@
where $\mathcal{P}_{\Gamma}^{\pm}=\sum_{\lambda:\mathrm{sign}(s_{\lambda})=\pm}\ket{\phi_{\lambda\Gamma}}\bra{\phi_{\lambda\Gamma}}$ |, so that $\mathcal{P}_{\Gamma}=\mathcal{P}_{\Gamma}^{+}+\mathcal{P}_{\Gamma}^{-}$. In particular, the covariant derivative is approximated by the forward finite difference formula in Eq. 21 and by the symmetric finite difference in Eq. 22. Finally, the PBC LSCM can be defined as:
@@
IV.2 Local PBC $\mathbb{Z}_{2}$ marker based on time-reversal symmetry
@@ V.1 Kane-Mele model
where $c_{i}^{\dagger}=(c_{i\uparrow}^{\dagger},c_{i\downarrow}^{\dagger})$ and sums on spin indices are implied, with the convention that if no spin matrices appear, they are contracted over the identity. In Equation V.1, $t$ is the nearest-neighbor hopping amplitude, and $\Delta$ is a staggered on-site potential, depending on the sublattice identified by $\tau_{i}\in\{0,1\}$. In the following, we will set $t=1$. The parameter $\lambda_{SO}$ is the intensity of the diagonal spin-orbit coupling, introduced as a complex hopping between second nearest-neighbors, where $\nu_{ij}=\mathrm{sign}(\mathbf{d}_{1}\times\mathbf{d}_{2})_{z}$ accounts for the direction of the hopping and $\mathbf{d}_{1,2}$ are unit vectors connecting the site $i$ to its second nearest-neighbor $j$. Finally, $\lambda_{R}$ is the amplitude of the Rashba term, which couples the two spin sectors and breaks $S_{z}$ symmetry. Here $\boldsymbol{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ is the vector of Pauli matrices, and ${\mathbf{e}}_{\langle ij\rangle}=\mathbf{d}_{\langle ij\rangle}\times\hat{\mathbf{z}}$ where $\mathbf{d}_{\langle ij\rangle}$ is the unit vector along the direction connecting site $i$ to site $j$.